OpenStudy (anonymous):

3e^2x=5e^4x solve for x

OpenStudy (espex):

If you take the ln of both sides, don't you en essence remove the e from your equation?

OpenStudy (espex):

Wolfram chose a different route:\[3e^2x=5e^4x \rightarrow 3e^2x-5e^4x=0 \rightarrow x(3e^2-5e^4)=0\] Divide both sides by 3e^2-5e^4 and you get x=0.